OpenStudy (lgbasallote):

another easy trig substitution please... \(\LARGE \int \frac{x^2}{(x^2 + a^2)^{\frac{3}{2}}} dx\)

OpenStudy (lgbasallote):

@Mimi_x3 and @apoorvk :D

OpenStudy (anonymous):

substitute\[x = a \tan \theta\]

OpenStudy (lgbasallote):

then...?

OpenStudy (mimi_x3):

then differntiate it..

OpenStudy (lgbasallote):

\(\large dx = a\sec^2 \theta\)

OpenStudy (lgbasallote):

\(d \theta\)

OpenStudy (mimi_x3):

\[\int\limits\frac{(atan\theta)^{2}}{((atan\theta)^{2}+a^{2})^{\frac{3}{2}}} *asec^{2}\theta \] then simply it..

OpenStudy (lgbasallote):

i cant see how it is gonna be simplified o.O

OpenStudy (lgbasallote):

ohwhait i do...

OpenStudy (lgbasallote):

the denom will be a^2 ( sec^2 theta)]^3/2 right?

OpenStudy (apoorvk):